OpenStudy (vera_ewing):

What exponential function represents the data in the table? x f(x) 3 64 4 256 5 1024 f(x) = x^4 f(x) = 4x f(x) = x^4 – 17 f(x) = 4x + 17

OpenStudy (vera_ewing):

I got f(x) = -x^4 + 17 but that's not one of the answer choices :(

OpenStudy (vera_ewing):

....no?

OpenStudy (dmndlife24):

I would just plug in the x for each answer choice and see if it gets its f(x)

OpenStudy (dmndlife24):

For example, f(x) = x^4 Plug in 4 for x... f(x) = 256

OpenStudy (vera_ewing):

So the answer is A?

OpenStudy (dmndlife24):

i think so

OpenStudy (zehanz):

Because 64=4x4x4, 256=4x4x4x4 and 1024=4x4x4x4x4, the function should be f(x)=4^x, but this is not one of your options.

OpenStudy (vera_ewing):

@nincompoop Is it A? Can you just make sure please?

OpenStudy (dmndlife24):

@ZeHanz thats what i was thinking

OpenStudy (nincompoop):

use QUALIFIED HELPER

OpenStudy (zehanz):

Maybe there's a "^"-character missing in answer b?

OpenStudy (zehanz):

Well, we KNOW the right answer f(x)=4^x, so imo you can move on to the next problem. No need to waste time on typesetting errors

OpenStudy (zehanz):

It's not one of your options, because all of them are wrong! Someone just made a typing error...